Gucchi:

Joy needs 8 feet of fabric for a project she is working on, but the store only sells the fabric in meters. One meter of fabric costs $1.15. How much will the fabric cost? [1 ft = 0.305m] $9.20 $2.12 $2.81 $30.16

Gucchi:

im confused on this one

Gucchi:

thinking a but dk

Gucchi:

@axie

Gucchi:

stop lying

Gucchi:

leave the question

Gucchi:

@joe348

Vocaloid:

Be careful with your units. You cannot multiply 8 * 1.15 because 8 is in feet and 1.15 is in dollars per meter. 1 foot = 0.305m So 8 feet = 0.305 * 8 = 2.44m Since you also know 1 m costs $1.15 it should be straightforward to calculate the cost of 2.44 m now

Gucchi:

wait so is it C? @vocaloid

Gucchi:

@snowflake0531

Gucchi:

helpppp

snowflake0531:

yes

snowflake0531:

first convert feet to meters then multiply it by the cost 8*0.305*1.15=2.806 which rounds to 2.81

Gucchi:

thanks